Status Report

NASA STS-117 FD08 Execute Package

By SpaceRef Editor
June 16, 2007
Filed under , , ,
NASA STS-117 FD08 Execute Package

Download complete document

FD7 MMT Crew Summary

The MMT met today and discussed mission progress, the Port OMS blanket repair, ISS attitude control anomalies, ascent hardware performance, and debris assessment results.

Port OMS Pod Blanket Repair:

The MMT received a status of the blanket repair verification testing and the corresponding schedule for completion. The blanket pull test has been completed which provides high confidence that the Saffil pins and staples will secure the blanket through the entry environment. The Arc Jet, Wind Tunnel, and Radiant Heat tests using the repaired configuration are now in work and most of these results are expected tomorrow. The Operations team reported that the EVA onboard training and procedure reviews for securing the blanket are in work.

ISS Attitude Control and Shuttle Consumables:

The MMT was briefed on the ISS attitude control issues and the latest troubleshooting efforts. The good news was that power from the US Segment was restored to the FGB and Soyuz which were operating on battery power. The attempts to regain the SM Central Computers were not successful today and as a result ISS continues to rely on CMG attitude control.

As a result of the ISS attitude control issues the MMT was also briefed on the Shuttle Cryo O2 and OMS/RCS consumables. The modified powerdown you performed today and delaying the O2 transfer will provide an additional docked day if required. Options to maximize RCS consumables are also in work which include the ability to handover attitude control from Shuttle to ISS CMG control, a modified ISS attitude while on verniers, or the use of Alt Dap. The details regarding these options are still in work.

Ascent Hardware Performance:

The SRB, ET, and SSME elements provided a post flight status to the MMT. Each element echoed that their system was nominal both from a hardware and performance perspective. The SSME AHMS controller worked as expected and the ET foam liberation was well within expected limits. An exceptionally clean ascent!

Imagery/Debris Assessment Team:

The RIB elevon TPS has been cleared. The team cleared this area by using imagery from the RPM pitch maneuver which was partially obscured by shadowing, SRB video of the elevon, and analysis. The analysis confirmed that even if TPS in this small region were damaged the result would be a tile turnaround issue only with no structural damage. The SRB video and RPM imagery indicate that there were no sig The team has no concerns with these tile.

SpaceRef staff editor.